# Time Makes Space: Emergence of Place Fields in Networks Encoding Temporally Continuous Sensory Experiences

### v1

@@ 3 Predictions
We also predict that the dimensionality of place fields is determined by the dimension in which temporally stable experiences change smoothly. As outlined in Sec. 2.3, place cells in our network prioritize encoding the most variable component within a localized region. The dimensionality of emergent place cells is therefore set by the projection of this encoded component onto the experience manifold (EM), and will be defined by the number of independent components of experience that change smoothly as an animal moves through a space—whether physical or abstract. To substantiate, we simulated an agent in 3D rooms and trained our RAE, similar to Sec. 2.2. We observed emergence of 3D place fields (Fig. 6b). This prediction is indirectly confirmed by Grieves et al. grievesPlacecellRepresentationVolumetric2020 , who observed 3D place cells in rodents navigating a lattice maze, a 3D environment that where rodents explore volumetrically. To test in physical spaces, VR setups with parametrically controlled sensory inputs could be employed. For abstract space, the prediction could be tested following procedures similar to the bird-deformation space used in Constantinescu et al. constantinescuOrganizingConceptualKnowledge2016 , or by placing rodents in 1D or 2D tracks while gradually changing auditory/olfactory signals within the enclosure or varying visual cues displayed on walls. We predict these continuous changes along the abstract dimension will extend previously generated place cells to higher dimensions.
@@ 4 Discussion
Acknowledgments: We thank Dori Derdikman, Genela Morris, and Shai Abramson for many illuminating discussions in the course of this work, which was supported by NIH CRCNS grant 1R01MH125544-01. We are also grateful to Pratik Chaudhari for his insightful suggestions for understanding the RAE dynamics.
@@ 1.2 Training Recurrent Autoencoder (RAE)
$\mathcal{L}=\frac{\lambda_{mse}}{D\cdot T\cdot B}\sum_{d,t,b}^{D,T,B}(\hat{e}_{d,t,b}-e_{d,t,b})^{2}+\frac{\lambda_{fr}}{N\cdot T\cdot B}\sum_{n,t,b}^{N,T,B}r_{n,t,b}^{2}$ (5)
@@
$\gamma=\frac{dt}{\tau}$ Decaying factor 10
@@
$T_{w}$ Width of the sampling window (see Fig. 1) 300 s
@@
$\Delta t$ The step size of how much to move the 1 s
sampling window forward

### v2

@@ 4 Predictions
We also predict that the dimensionality of place fields is determined by the dimension in which temporally stable experiences change smoothly. As outlined in Sec. 3.2, place cells in our network prioritize encoding the most variable component within a localized region. The dimensionality of emergent place cells is therefore set by the projection of this encoded component onto the experience manifold (EM), and will be defined by the number of independent components of experience that change smoothly as an animal moves through a space—whether physical or abstract. To substantiate, we simulated an agent in 3D rooms and trained our RAE, similar to Sec. 3.1. We observed the emergence of 3D place fields (Fig. 6b). This prediction is indirectly confirmed by Grieves et al. grievesPlacecellRepresentationVolumetric2020 , who observed 3D place cells in rodents navigating a lattice maze, a 3D environment that where rodents explore volumetrically. To test in physical spaces, VR setups with parametrically controlled sensory inputs could be employed. For abstract space, the prediction could be tested following procedures similar to the bird-deformation space used in Constantinescu et al. constantinescuOrganizingConceptualKnowledge2016 , or by placing rodents in 1D or 2D tracks while gradually changing auditory/olfactory signals within the enclosure or varying visual cues displayed on walls. We predict these continuous changes along the abstract dimension will extend previously generated place cells to higher dimensions.
@@ 5 Discussion
Acknowledgments: We thank Dori Derdikman, Genela Morris, and Shai Abramson for many illuminating discussions in the course of this work, which was supported by NIH CRCNS grant 1R01MH125544-01 and in part by the NSF and DoD OUSD (R&E) under Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence). VB was supported in part by the Eastman Professorship at Balliol College, Oxford. We are also grateful to Pratik Chaudhari for his insightful suggestions on designing the RAE and analyzing its dynamics.
@@ 1.2 Training Recurrent Autoencoder (RAE)
$\mathcal{L}=\frac{\lambda_{mse}}{D\cdot T\cdot B}\sum_{d,t,b}^{D,T,B}(\hat{e}_{d,t,b}-e_{d,t,b})^{2}+\frac{\lambda_{fr}}{N}\sum_{n}^{N}\left(\sum_{t,b}^{T,B}r_{n,t,b}\right)^{2}$ (5)
@@
$\gamma=\frac{dt}{\tau}$ Decaying factor 0.1
@@
$T_{w}$ Width of the sampling window (see Fig. 1 caption) 300 s
@@
$\Delta t$ The step size of how much to move the sampling 1 s
window forward